OpenStudy (anonymous):

Simplify+ find the domain? root(3x-5) * root(3x-5) - (2x+1)

hartnn (hartnn):

could you simplify ?

OpenStudy (anonymous):

how would you do that

OpenStudy (anonymous):

\[\sqrt{3x-5} \times \sqrt{3x-5} - (2x+1)\]

hartnn (hartnn):

use this , \(\sqrt{stuff}\sqrt{stuff}=stuff\)

hartnn (hartnn):

so, \(\sqrt{3x-5} \times \sqrt{3x-5} =...?\)

OpenStudy (anonymous):

ohhh so its 3x-5 - 2x+1

hartnn (hartnn):

3x-5 - (2x+1) actually so, 3x-5 - 2x-1 =... ?

OpenStudy (anonymous):

x-6!

OpenStudy (anonymous):

Oh I did not know the two roots makes zero oh x_x

OpenStudy (anonymous):

thnx!

hartnn (hartnn):

what about the domain ?

hartnn (hartnn):

and 2 roots doesn't make a "zero" actually , \(\large \sqrt x\sqrt x = x^{0.5}x^{0.5}=x^{0.5+0.5}=x^1=x\)

OpenStudy (anonymous):

ohhh I see thanks

hartnn (hartnn):

and domain ?

OpenStudy (anonymous):

x is greater than or equal to 3/5?

hartnn (hartnn):

3/5 or 5/3 ?

hartnn (hartnn):

\(3x-5\ge 0 \\ 3x \ge 5 \\x \ge 5/3\)

OpenStudy (anonymous):

oh sorry ment 5/3

hartnn (hartnn):

then you are correct! good work! :)

OpenStudy (anonymous):

thnx!!!

hartnn (hartnn):

welcome ^_^